# Multi-Agent Guided Policy Optimization

### v1

@@ 4 Method
We begin by presenting the theoretical formulation and guarantee of monotonic policy improvement in the tabular setting. For clarity, we initially assume full observability—i.e., all agents observe the global state $s$, reducing the setting to a cooperative Markov game. We will return to the partially observable case in the following implementation section.
@@ 4.1 Multi-Agent Guided Policy Optimization
Our algorithm maintains a centralized guider policy with an autoregressive structure over agent actions: $\bm{\mu}(\bm{a}|s)=\mu^{i_{1}}(a^{i_{1}}|s)\mu^{i_{2}}(a^{i_{2}}|s,a^{i_{1}})\dots\mu^{i_{n}}(a^{i_{n}}|s,\bm{a}^{i_{1:n-1}})$, where $i_{1:m}$ (with $m\leq n$) denotes an ordered subset $\{i_{1},...,i_{m}\}$ of the agent set $\mathcal{N}$, specifying the execution order. The decentralized learner policy is defined as: $\bm{\pi}(\bm{a}|s)=\prod_{j=1}^{n}\pi^{i_{j}}(a^{i_{j}}|s)$. for any ordering $i_{1:n}$, implying that all agents act independently.
Building on this structure, MAGPO optimizes the centralized guider and decentralized learner policies through an iterative four-step procedure inspired by the GPO framework:
@@
The first step allows MAGPO to perform coordinated exploration using a joint policy. In the second step, the guider is updated using the Policy Mirror Descent (PMD) framework [45], which solves the following optimization:
@@
where $\eta_{k}$ is the learning rate. PMD is a general policy gradient method that encompasses popular algorithms like PPO and SAC. Here, we use PMD for theoretical convenience and instantiate it with PPO-style updates in the practical implementation section. In the final step, we perform guider backtracking, where the guider is reset to the current learner policy. Theoretically, this is always feasible since any decentralized policy $\bm{\pi}$ defines a valid autoregressive joint policy $\bm{\mu}$ by simply ignoring the conditioning on past actions.
Based on the framework introduced above, we can establish the monotonic improvement guarantee for MAGPO.
@@
To further clarify the structure of MAGPO, we show that its learner updates can be interpreted as sequential advantage-based updates—a procedure known to ensure monotonic improvement in multi-agent settings [19]. We begin with the following lemma:
Lemma 1 (Multi-Agent Advantage Decomposition [19]).
@@
This shows that MAGPO’s learner updates are equivalent to performing sequential advantage-weighted policy updates. Importantly, unlike methods such as HARL which update agents one at a time, MAGPO allows for simultaneous updates of all agent policies. This enables parallel execution and improves scalability to large agent populations. Moreover, HARL requires heterogeneous agents to guarantee policy improvement, while MAGPO works with either homogeneous or heterogeneous agents, allowing it to benefit from parameter sharing—a widely adopted practice that significantly improves efficiency and generalization in MARL [14, 37, 6].
@@ 4.2 Practical Implementation
In this subsection, we describe the practical implementation of MAGPO. Our implementation is based on the original GPO-Clip framework, extended to the multi-agent setting. The key difference is that the guider in MAGPO is a sequential execution policy. Since MAGPO is compatible with any autoregressive CTCE method, we do not specify the exact encoder, decoder, or attention mechanisms used. Instead, we present general training objectives for both the guider and learner components.
@@
$\mathcal{L}(\phi)=-\frac{1}{Tn}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\Big{[}\min\Big{(}r_{t}^{i_{j}}(\phi)\hat{A}_{t},$ $\text{clip}(r_{t}^{i_{j}}(\phi),\epsilon,\delta)\hat{A}_{t}\Big{)}$ (8)
$-m_{t}^{i_{j}}\text{D}_{\text{KL}}\big{(}\mu^{i_{j}}_{\phi}(\cdot|s_{t},\bm{a}_{t}^{i_{1:j-1}}),\pi_{\theta}^{i_{j}}(\cdot|o_{t}^{i_{j}})\big{)}\Big{]},$
@@
$\text{clip}(r_{t}^{i_{j}}(\phi),\epsilon,\delta)=\text{clip}\left(\text{clip}\left(\frac{\mu_{\phi}^{i_{j}}(a^{i_{j}}_{t}|s_{t},\bm{a}_{t}^{i_{1:j-1}})}{\pi_{\theta}^{i_{j}}(a^{i_{j}}_{t}|o_{t}^{i_{j}})},\frac{1}{\delta},\delta\right)\frac{\pi_{\theta}^{i_{j}}(a^{i_{j}}_{t}|o_{t}^{i_{j}})}{\mu_{\phi_{\text{old}}}^{i_{j}}(a^{i_{j}}_{t}|s_{t},\bm{a}_{t}^{i_{1:j-1}})},1-\epsilon,1+\epsilon\right).$ (9)
This objective has two modifications compared to the standard one: a double clipping function clip$(\cdot,\epsilon,\delta)$ and a mask function $m^{i_{j}}(\delta)$, both controlled by a new hyperparameter $\delta>1$ , which bounds the ratio between guider and learner policies within $(\frac{1}{\delta},\delta)$. The inner clip in the double clipping function stops the gradient when the advantage signal encourages the guider to drift too far from the learner. The mask function ensures the KL loss is only applied when this ratio constraint is violated. The advantage estimate $\hat{A}_{t}$ is computed via generalized advantage estimation (GAE) [32] with value functions.
@@
$\mathcal{L}(\theta)=\frac{1}{Tn}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\Big{[}\text{D}_{\text{KL}}\big{(}\pi_{\theta}^{i_{j}}(\cdot|o_{t}^{i_{j}}),$ $\mu^{i_{j}}_{\phi}(\cdot|s_{t},\bm{a}_{t}^{i_{1:j-1}})\big{)}$ (10)
$-\lambda\min\left(r_{t}^{i_{j}}(\theta)\hat{A}_{t},\text{clip}(r_{t}^{i_{j}}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\Big{]},$
@@
The auxiliary RL objective helps maximize the utility of collected trajectories. Since the behavior policy (guider) is kept close to the learner, this term approximates an on-policy objective. In principle, we could apply sequential updates to each individual policy—analogous to HAPPO—to preserve the theoretical guarantees of monotonic improvement. However, this makes the learner updates non-parallelizable and incompatible with parameter sharing. To ensure efficiency and scalability, we instead adopt a MAPPO-style update: all learners are updated jointly and in parallel. The auxiliary RL term can be treated as optional and controlled by $\lambda$.
@@ 5 Experiments
We evaluate MAGPO by comparing it against several SOTA baseline algorithms from the literature. Specifically, we consider two CTCE methods—Sable [23] and MAT [44]—two standard baseliens—MAPPO [47] and IPPO [10]—and a vanilla implementation of on-policy CTDS, which can be viewed as MAGPO without double clipping, masking, and the RL auxiliary loss. For the joint policy in both MAGPO and CTDS, we use Sable as the default backbone. All algorithms are implemented using the JAX-based MARL library Mava [9].
Evaluation protocol.
We follow the evaluation protocol from Mahjoub et al. [23]. Each algorithm is trained with 10 independent seeds per task. Training is conducted for 20 million environment steps, with 122 evenly spaced evaluation checkpoints. At each checkpoint, we record the mean episode return over 32 evaluation episodes, along with any task-specific metrics (e.g., win rate). For task-level results, we report the mean and 95% confidence intervals. For aggregate performance across entire environment suites, we report the min-max normalized interquartile mean (IQM) with 95% stratified bootstrap confidence intervals.
@@
We evaluate MAGPO on a diverse suite of JAX-based multi-agent benchmarks, including 3 tasks in CoordSum (introduced in this paper), 15 tasks in Robotic Warehouse (RWARE) [27], 7 tasks in Level-based foraging (LBF) [7], 4 tasks in Connector [3], 11 tasks in The StarCraft Multi-Agent Challenge in JAX (SMAX) [31], and 3 tasks in the Multi-agent Particle Environment (MPE) [22]. The CoordSum environment, introduced in this paper, reflects the didactic examples discussed in Section 3, where agents must coordinate to output integers that sum to a given target without using fixed strategies. A detailed description is provided in Appendix B.
Hyperparameters.
Most of the baseline results are taken directly from Mava, which provides per-task tuned hyperparameters. For MAGPO, we adopt the hyperparameters of the corresponding CTCE method (Sable or MAT) and tune only the additional hyperparameters $\delta$ and $\lambda$ within a small search space. For the CoordSum environment, which is newly introduced, we tune all hyperparameters from scratch for each algorithm. Full details are provided in Appendix C.3.
Figure 2: The sample efficiency curves aggregated per environment suite. For each environment, results are aggregated over all tasks and the min–max normalized inter-quartile mean with 95% stratified bootstrap confidence intervals are shown.
Figure 3: The overall aggregated probability of improvement for MAGPO compared to other baselines for that specific environment. A score of more than 0.5 where confidence intervals are also greater than 0.5 indicates statistically significant improvement over a baseline for a given environment [1].
@@ 5.1 Main Results
Figure 2 presents the per-environment aggregated sample-efficiency curves. Our results show that MAGPO achieves state-of-the-art performance across all CTDE methods and even outperforms CTCE methods on a subset of tasks. Specifically, MAGPO surpasses all CTDE baselines on 33 out of 43 tasks, and outperforms all baselines on 19 out of 43 tasks. Figure 3 reports the probability of improvement of MAGPO over other baselines. MAGPO emerges as the most competitive CTDE method and performs comparably to the SOTA CTCE method Sable in three benchmark environments. Comparing MAGPO to CTDS reveals a significant performance gap in the CoordSum and RWARE domains, suggesting that in these environments, the CTCE teacher may learn policies that are not decentralizable—rendering direct policy distillation ineffective. Additional tabular results and environment/task-level aggregation plots are provided in Appendix C.2.
@@ 5.2 Ablations and Discussions
MAGPO’s performance intuitively depends heavily on the capability of the guider, which corresponds to the performance of the underlying CTCE method. In Figure 4(a), we evaluate MAGPO on two tasks where Sable and MAT exhibit different behaviors. In simple_spread_10ag, MAT performs significantly worse, resulting in poor performance of MAGPO when using MAT as the guider. In contrast, on large-8ag, MAT outperforms Sable, leading to better performance of MAGPO with MAT. This dependency could be seen as a limitation, but we argue it serves as a feature: MAGPO effectively bridges CTCE and CTDE. In many practical applications that require decentralized policies, MAGPO enables advances in CTCE methods to directly benefit CTDE methods as well—facilitating the co-development of both paradigms.
@@
MAGPO introduces a hyperparameter $\delta$ to regulate the guider’s deviation from the learner. A smaller $\delta$ enforces a stricter constraint, keeping the guider closer to the learner policy; a larger $\delta$ allows the guider more freedom, potentially enabling it to explore regions of the policy space that are difficult or even unreachable under decentralized constraints. In Figure 4(b), we assess MAGPO’s performance under varying $\delta$ values on two tasks. In CoordSum-5x20, a smaller $\delta$ yields better performance because the centralized guider tends to learn an undecentralizable policy, which must be restricted to improve imitability. Conversely, in medium-4ag-hard, the guider policy is more directly imitable, and restricting it too tightly hinders learning. These observations show the importance of tuning $\delta$ based on the task’s structure and imitation feasibility.
@@
MAGPO incorporates an RL auxiliary loss in the learner update to better utilize collected data and stabilize learning. This component is critical to MAGPO’s ability to match or even outperform the corresponding CTCE methods. Without this auxiliary loss, the guider—when compared to the CTCE method—acts as a more conservative version that may converge more slowly. However, as shown in Figure 5(a), a properly tuned $\lambda$ can significantly improve performance. Conceptually, if the guider updates toward an undecentralizable direction and the learner pulls it back (due to the imitation constraint), and the learner itself does not learn through RL, then this back-and-forth may repeatedly stall progress. By incorporating RL updates, the learner can “counter-supervise” the guider, helping it discover more decentralizable update directions. Furthermore, in Figure 5(b), we test applying the same RL auxiliary loss to a CTDS method. The results show limited benefit. This is because in CTDS, the behavioral policy is the teacher, and there is no constraint enforcing alignment with the student policy. If the teacher-student gap is too large, the on-policy RL loss on the student provides little benefit.
@@
While much of our analysis has focused on asymmetries in the policy and action spaces, observation asymmetry is equally critical. This asymmetry can significantly hinder imitation performance, as the individual (decentralized) policy is expected to mimic decisions made under a richer observation space—decisions that may be unrealizable or uninterpretable given only local observations. Currently, CTCE methods such as MAT and Sable condition on the union of agents’ partial observations, whereas individual policies are limited to their own local views. This mismatch creates an imitation gap, even when the underlying joint policy is theoretically decentralizable. CTDS methods can also suffer from this gap due to the same asymmetry. MAGPO addresses this issue by controlling the divergence between the guider and the learner through the parameter $\delta$. In scenarios with significant observation asymmetry, using a smaller $\delta$ encourages the guider to remain closer to what is achievable by partial observable decentralized policies, thereby improving imitation performance. Moreover, privileged information—beyond the union of partial observations—is often available during centralized training (e.g., the true global state), although we do not explore it in this paper. Providing such privileged signals to the guider could further enhance its ability to supervise decentralized policies under partial observability. This idea aligns with the GPO framework in the single-agent setting [20]. Extending this principle to MAGPO offers a promising direction for future work, especially in practical domains like robotics, where centralized training with full state access is common, but execution must rely on partial and noisy local observations.

### v2

@@ 4 Method
We begin by presenting the theoretical formulation and guarantee of monotonic policy improvement in the tabular setting. For simplicity, we initially assume full observability—i.e., all agents observe the global state $s$, reducing the setting to a cooperative Markov game (Littman, 1994). We will return to the partially observable case in the subsequent implementation section.
@@ 4.1 Multi-Agent Guided Policy Optimization
Our algorithm maintains a centralized guider policy with an autoregressive structure over agent actions: $\bm{\mu}(\bm{a}|s)=\mu^{i_{1}}(a^{i_{1}}|s)\mu^{i_{2}}(a^{i_{2}}|s,a^{i_{1}})\dots\mu^{i_{n}}(a^{i_{n}}|s,\bm{a}^{i_{1:n-1}})$, where $i_{1:m}$ (with $m\leq n$) denotes an ordered subset $\{i_{1},...,i_{m}\}$ of the agent set $\mathcal{N}$, specifying the execution order. The decentralized learner policy is defined as: $\bm{\pi}(\bm{a}|s)=\prod_{j=1}^{n}\pi^{i_{j}}(a^{i_{j}}|s)$ for any ordering $i_{1:n}$, implying that all agents act independently.
Building on this structure, MAGPO optimizes the centralized guider and decentralized learner policies through an iterative four-step procedure inspired by the GPO framework (Li et al., 2025):
@@
The first step allows MAGPO to perform coordinated exploration using a joint policy. In the second step, the guider is updated using the Policy Mirror Descent (PMD) framework (Xiao, 2022), which solves the following optimization:
@@
where $Q^{\bm{\mu}_{k}}$ is the Q-function of guider and $\eta_{k}$ is the learning rate. As discussed in Section 2.1, PMD is a general policy gradient framework that subsumes popular algorithms such as PPO and TRPO. Here, we adopt PMD for theoretical clarity and instantiate it using PPO-style updates in our practical implementation. Additional details are provided in Appendix A. In the final step, we perform guider backtracking, where the guider is reset to the current learner policy. Theoretically, this is always feasible since any decentralized policy $\bm{\pi}$ defines a valid autoregressive joint policy $\bm{\mu}$ by simply ignoring the conditioning on past actions.
Based on the framework introduced above, we have the following theorem for MAGPO.
@@
To further clarify the structure of MAGPO, we show that its learner updates can be interpreted as sequential advantage-based updates—a procedure known to ensure monotonic improvement in multi-agent settings (Kuba et al., 2022). We begin with the following lemma:
Lemma 1 (Multi-Agent Advantage Decomposition (Kuba et al., 2022)).
@@
This shows that MAGPO’s learner updates are equivalent to performing sequential advantage-weighted policy updates. Importantly, unlike methods such as HARL which update agents one at a time, MAGPO allows for simultaneous updates of all agent policies. This enables parallel training and improves scalability to large agent populations. Moreover, HARL requires heterogeneous agents to guarantee policy improvement, while MAGPO works with either homogeneous or heterogeneous agents, allowing it to benefit from parameter sharing—a widely adopted practice that significantly improves efficiency and generalization in MARL (Gupta et al., 2017; Terry et al., 2020; Christianos et al., 2021a).
@@ 4.2 Practical Implementation
In this subsection, we describe the practical implementation of MAGPO. Our implementation is based on the original GPO-clip framework (Li et al., 2025), extended to the multi-agent setting. The key difference is that the guider in MAGPO is a sequential execution policy. Since MAGPO is compatible with any autoregressive CTCE method, we do not specify the exact encoder, decoder, or attention mechanisms used. Instead, we present general training objectives for both the guider and learner components.
@@
$\mathcal{L}(\phi)=-\frac{1}{Tn}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\Big[\min\Big(r_{t}^{i_{j}}(\phi)\hat{A}_{t},\text{clip}($ $r_{t}^{i_{j}}(\phi),\epsilon,\delta)\hat{A}_{t}\Big)$- (9)
$m_{t}^{i_{j}}(\delta)\text{D}_{\text{KL}}\big(\mu^{i_{j}}_{\phi}(\cdot|s_{t},\bm{a}_{t}^{i_{1:j-1}}),\pi_{\theta}^{i_{j}}(\cdot|o_{t}^{i_{j}})\big)\Big],$
@@
$\text{clip}(r_{t}^{i_{j}}(\phi),\epsilon,\delta)=\text{clip}\left(\text{clip}\left(\frac{\mu_{\phi}^{i_{j}}(a^{i_{j}}_{t}|s_{t},\bm{a}_{t}^{i_{1:j-1}})}{\pi_{\theta}^{i_{j}}(a^{i_{j}}_{t}|o_{t}^{i_{j}})},\frac{1}{\delta},\delta\right)\frac{\pi_{\theta}^{i_{j}}(a^{i_{j}}_{t}|o_{t}^{i_{j}})}{\mu_{\phi_{\text{old}}}^{i_{j}}(a^{i_{j}}_{t}|s_{t},\bm{a}_{t}^{i_{1:j-1}})},1-\epsilon,1+\epsilon\right).$
[…]
@@
$\mathcal{L}(\theta)=\frac{1}{Tn}\sum_{j=1}^{n}\sum_{t=0}^{T-1}\Big[\text{D}_{\text{KL}}\big(\pi_{\theta}^{i_{j}}(\cdot|o_{t}^{i_{j}}),$ $\mu^{i_{j}}_{\phi}(\cdot|s_{t},\bm{a}_{t}^{i_{1:j-1}})\big)$- (10)
$\lambda\min\left(r_{t}^{i_{j}}(\theta)\hat{A}_{t},\text{clip}(r_{t}^{i_{j}}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\Big],$
@@
The auxiliary RL objective helps maximize the utility of collected trajectories. Since the behavior policy (guider) is kept close to the learner, this term approximates an on-policy objective. In principle, we could apply sequential updates to each individual policy—analogous to HAPPO—to preserve the theoretical guarantees of monotonic improvement. However, this makes the learner updates non-parallelizable and incompatible with parameter sharing. Since no performance benefits are observed from using HAPPO, we instead adopt a MAPPO-style update: all learners share parameters and are updated jointly and in parallel. The auxiliary RL term can be treated as optional and controlled by $\lambda$.
((a))
((a))
((b))
((c))
((d))
((e))
((f))
Figure 2: The sample efficiency curves aggregated per environment suite, where dashed lines represent the CTCE methods. For each environment, results are aggregated over all tasks and the min–max normalized inter-quartile mean with 95% stratified bootstrap confidence intervals are shown.
Figure 3: The overall aggregated probability of improvement for MAGPO compared to other baselines for that specific environment. A score of more than 0.5 where confidence intervals are also greater than 0.5 indicates statistically significant improvement over a baseline for a given environment (Agarwal et al., 2021).
@@ 5 Experiments
We evaluate MAGPO by comparing it against several SOTA baseline algorithms from the literature. Specifically, we consider two CTCE methods—Sable (Mahjoub et al., 2025) and MAT (Wen et al., 2022)—two CTDE baselines—MAPPO (Yu et al., 2022) and HAPPO (Kuba et al., 2022)—and a vanilla implementation of on-policy CTDS, which can be viewed as MAGPO without double clipping, masking, and the RL auxiliary loss. For the joint policy in both MAGPO and CTDS, we use Sable as the default backbone. All algorithms are implemented using the JAX-based MARL library Mava (de Kock et al., 2023).
Evaluation & Hyperparameters.
We follow the evaluation protocol from Mahjoub et al. (2025). Each algorithm is trained with 10 independent seeds per task. Training is conducted for 20 million environment steps, with 122 evenly spaced evaluation checkpoints. At each checkpoint, we record the task-specific metrics (e.g., mean episode return and win rate) over 32 evaluation episodes. For task-level results, we report the mean and 95% confidence intervals. For aggregate performance across entire environment suites, we report the min-max normalized interquartile mean (IQM) with 95% stratified bootstrap confidence intervals. The hyperparameters are tuned on each task for each algorithm, which are detailed in Appendix C.3.
@@
We evaluate MAGPO on a diverse suite of JAX-based multi-agent benchmarks, including 4 tasks in CoordSum (introduced in this paper), 7 tasks in Level-based foraging (LBF) (Christianos et al., 2021b), 4 tasks in Connector (Bonnet et al., 2024), 3 tasks in the Multi-agent Particle Environment (MPE) (Lowe et al., 2017), 15 tasks in Robotic Warehouse (RWARE) (Papoudakis et al., 2020), and 10 tasks in The StarCraft Multi-Agent Challenge in JAX (SMAX) (Rutherford et al., 2024). The CoordSum environment, introduced in this paper, reflects the didactic examples discussed in Section 3, where agents must coordinate to output integers that sum to a given target without using fixed strategies. A detailed description is provided in Appendix B.
@@ 5.1 Main Results
Figure 2 presents the per-environment aggregated sample-efficiency curves. Our results show that MAGPO achieves state-of-the-art performance across all CTDE methods and even outperforms CTCE methods on a subset of tasks. Specifically, MAGPO surpasses all CTDE baselines on 32 out of 43 tasks, and outperforms all baselines on 20 out of 43 tasks. Figure 3 reports the probability of improvement of MAGPO over other baselines. MAGPO emerges as the most competitive CTDE method and performs comparably to the SOTA CTCE method Sable in three benchmark environments. Comparing MAGPO to CTDS reveals a significant performance gap in the CoordSum and RWARE domains, suggesting that in these environments, the CTCE teacher may learn policies that are not decentralizable—rendering direct policy distillation ineffective. Additional tabular results and environment/task-level aggregation plots are provided in Appendix C.2.
((a))
((b))
@@ 5.2 Ablations and Discussions
MAGPO’s performance intuitively depends heavily on the capability of the guider, which corresponds to the performance of the underlying CTCE method. In Figure 4(a), we evaluate MAGPO on two tasks where Sable and MAT exhibit different performance. In simple_spread_10ag, MAT performs significantly worse, resulting in poor performance of MAGPO when using MAT as the guider. In contrast, on large-8ag, MAT outperforms Sable, leading to better performance of MAGPO with MAT. This dependency could be seen as a limitation, but it actuall serves as a core feature: MAGPO effectively bridges CTCE and CTDE. In many practical applications that require decentralized policies, MAGPO enables advances in CTCE methods to directly benefit CTDE methods as well—facilitating the co-development of both paradigms.
@@
MAGPO introduces a hyperparameter $\delta$ to regulate the guider’s deviation from the learner, which most strongly influences its performance. A smaller $\delta$ enforces a stricter constraint, keeping the guider closer to the learner policy; a larger $\delta$ allows the guider more freedom, potentially enabling it to explore regions of the policy space that are difficult or even unreachable under decentralized constraints. In Figure 4(b), we assess MAGPO’s performance under varying $\delta$ values on two tasks. In CoordSum-5x20-80, a smaller $\delta$ yields better performance because the centralized guider tends to learn a policy that is not decentralizable, which must be restricted to improve imitability. Conversely, in medium-4ag-hard, the guider policy is more directly imitable, and restricting it too tightly hinders learning. These observations show the importance of tuning $\delta$ based on the task’s structure and imitation feasibility.
@@
MAGPO incorporates an RL auxiliary loss in the learner update to better utilize collected data and stabilize learning. This component is not as important as the $\delta$, but a properly tuned $\lambda$ can also improve performance, as shown in Figure 5(a). To understand this, consider the guider’s RL objective is towards an undecentralizable direction and the learner pulls it backward (due to the imitation constraint), then this back-and-forth may repeatedly stall progress. By incorporating RL updates, the learner can “counter-supervise” the guider, helping it discover more decentralizable update directions. Furthermore, in Figure 5(b), we test applying the same RL auxiliary loss to a CTDS method. The results show limited benefit. This is because in CTDS, the behavioral policy is the teacher, which is not enforced to align with the student. If the teacher-student gap is too large, the collected data is off-policy, thus on-policy RL loss on the student provides little benefit.
@@
While most of our analysis has focused on asymmetries in the policy and action spaces, observation asymmetry is equally critical. In the current framework, CTCE methods like MAT and Sable condition on the union of agents’ partial observations, whereas individual policies are limited to their own local views. This mismatch creates an imitation gap, making direct imitation methods like CTDS fail, even when the underlying joint policy is decentralizable. MAGPO addresses this issue similar to the single-agent setting (Li et al., 2025), by controlling the divergence between the guider and the learner through the parameter $\delta$. In addition, privileged information—beyond the union of partial observations—is often available during centralized training (e.g., the true global state), although we do not explore it in this paper. Providing such privileged signals to the guider could further enhance its ability to supervise decentralized policies under partial observability.
Figure 6: Effect of model capacity on performance.
Model capacity.
In addition to policy and observation asymmetry, asymmetry can also arise from mismatched model capacity between training and deployment. In many practical systems, a high-capacity centralized model or teacher is used during training, while a compact decentralized policy must be deployed due to compute or latency constraints (e.g., large LLMs distilled for real-time inference, or vision-language-action models paired with lightweight controllers for high-frequency control). MAGPO naturally accommodates this setting by explicitly constraining the centralized guider to remain imitable by decentralized actors throughout training. To evaluate this property, we conduct an additional experiment on smacv2_5_units, where all training-time networks (e.g., critics and centralized policies) are kept at full capacity, while only the hidden dimension of the deployable decentralized actors is reduced at evaluation time. This setup simulates a realistic distillation scenario in which a large teacher must be compressed for deployment. As shown in Figure 6, MAGPO consistently outperforms CTDS across all evaluated actor capacities, and the performance gap widens as the deployed actor becomes smaller. Compared to MAPPO, both MAGPO and CTDS—being teacher–student frameworks—transfer knowledge from large models more effectively than value-based supervision alone. Importantly, MAGPO degrades more gracefully under severe capacity constraints, indicating that constraining the teacher to remain aligned with decentralized realizability improves robustness to deployment compression.